# Time Series Modeling for Heart Rate Prediction: From ARIMA to Transformers

### v2

@@ I Introduction
Cardiovascular disease is one of the top causes of death globally, claiming millions of lives annually. To prevent any unforeseen events during surgery or the post-operative recovery, it is imperative to create forecasting models that allow healthcare facilities to monitor patients’ vital bio-metric signs, such as heart rate, blood pressure, and electrocardiogram (ECG) [1, 2, 3]. Various statistical and machine learning models, including ARIMA and Prophet, are frequently employed in these contexts due to their proficiency in handling time series data. However, these traditional models exhibit significant limitations that undermine their reliability and performance, particularly in managing the complexity and non-linearity inherent in cardiovascular conditions. Consequently, given the critical need for effective forecasting models in cardiovascular disease management and the inadequacies of traditional models, there is an urgent necessity for more advanced approaches.
@@
Recognizing the limitations of traditional models in handling large datasets with extended time windows, deep learning-based methods have gained popularity in forecasting. For example, Zheng et al. [5] combined the Synthetic Minority Over-sampling Technique with XGBoost for fraud detection, improving classification efficiency. Zhou et al. [6] demonstrated how combining reinforcement learning with Generalized Second-Price auctions can significantly improve model accuracy and robustness in ad ranking, illustrating its application in advertising. Furthermore, AI is being leveraged in numerous other fields, including image processing for better resolution [7, 8, 9], RNN from adder’s perspective [10], algorithm for portfolio optimization [11], showcasing its versatility and effectiveness in diverse applications. However, these methods have not been applied to cardiovascular diseases, which present unique complexities. To address these challenges, more advanced models capable of handling the complexity and non-linearity of cardiovascular disease data are needed. Deep learning models, such as neural networks [12, 13, 14, 15, 16, 17], convolutional neural networks (CNNs) [18, 19, 20, 21, 22, 23], recurrent neural networks (RNNs), and transformers [24, 25, 26, 27, 28, 29, 17, 30, 31, 32], are well-suited for processing large, complex datasets and offer a promising solution for improving cardiovascular disease forecasting.
@@ II-A Data Description and Preparation
The heart rate time series dataset used in this study was obtained from the MIT-BIH Database. This dataset contained heart rate measurements for four different individuals. For T1 and T2, each series included 1800 equally spaced instant heart rate readings (in beats per minute) obtained at intervals of 0.5 seconds. T3 and T4 were obtained in the same way with 950 and 851 measurements each. Z-score analysis was utilized to identify potential outliers in the dataset. After a thorough investigation, it was determined that these outliers were not due to input errors; therefore, they were retained for further analysis. Min-max normalization was applied to all models to ensure that all values fall within the range of 0 to 1 and to standardize the input data scale across each model. This step is crucial when working with deep learning models because it facilitates faster convergence during training and enhances overall model performance. Normalizing the data helps mitigate the effects of varying feature scales, leading to improved numerical stability and more reliable results.
@@ II-B Framework
For each of the four-time series datasets (T1, T2, T3, T4), individual models were constructed using Seasonal-ARIMA, Prophet, LSTM, TSMixerx, TimesNet, TCN, PatchTST, and iTransformer. Drawing inspiration from [5], we recognized the necessity to address the complexities in optimizing our models, particularly focusing on improving model accuracy and robustness.
@@ II-C Time Series Models Training
Prophet: We utilized the Facebook Prophet model as another benchmark. This model was employed with parameters such as changepoint_prior_scale, seasonality_prior_scale, and n_changepoints tuned through cross-validation. We also combined with adjustments from [6] aimed to optimize forecasting accuracy by controlling the model’s sensitivity to trend changes and seasonal fluctuations.
@@
Long Short-Term Memory (LSTM) Network [33, 34, 35, 9]: The LSTM model included two LSTM layers followed by a dense layer. Heart rate data was normalized using MinMaxScaler. We use three layers of LSTM with hidden dimension of 32. The model was trained using the MSE loss function to enhance performance.
Temporal Convolutional Network (TCN) has been used in time series prediction with huge success [36, 37, 38, 39, 40]. We trained the model using dilated temporal convolutions to capture long-range dependencies in the heart rate data. Five layers of dilated convolutions with residual connections were employed to enhance training efficiency. The heart rate data was normalized before training. We set filter size to be three and dilation factor to be two. The MSE loss function used to guide the training process, ensuring effective learning of temporal patterns. We followed the work of Zhou et al. [6] to enhance model accuracy and robustness by dynamically balancing multiple input factors.

### v3

@@ I Introduction
Cardiovascular disease is one of the top causes of death globally, claiming millions of lives annually. To prevent any unforeseen events during surgery or the post-operative recovery, it is imperative to create forecasting models that allow healthcare facilities to monitor patients’ vital bio-metric signs, such as heart rate, blood pressure, and electrocardiogram (ECG). Various statistical and machine learning models, including ARIMA and Prophet, are frequently employed in these contexts due to their proficiency in handling time series data. However, these traditional models exhibit significant limitations that undermine their reliability and performance, particularly in managing the complexity and non-linearity inherent in cardiovascular conditions. Consequently, given the critical need for effective forecasting models in cardiovascular disease management and the inadequacies of traditional models, there is an urgent necessity for more advanced approaches.
@@
Recognizing the limitations of traditional models in handling large datasets with extended time windows, deep learning-based methods have gained popularity in forecasting. AI is being leveraged in numerous other fields, including image processing for better resolution, RNN from adder’s perspective, algorithm for portfolio optimization, showcasing its versatility and effectiveness in diverse applications. However, these methods have not been applied to cardiovascular diseases, which present unique complexities. To address these challenges, more advanced models capable of handling the complexity and non-linearity of cardiovascular disease data are needed. Deep learning models, such as neural networks [2, 3, 4], convolutional neural networks (CNNs), recurrent neural networks (RNNs), and transformers [5, 6, 7, 8], are well-suited for processing large, complex datasets and offer a promising solution for improving cardiovascular disease forecasting.
@@ II-A Data Description and Preparation
The heart rate time series dataset used in this study was obtained from the MIT-BIH Database. This dataset contained heart rate measurements for four different individuals. For T1 and T2, each series included 1800 equally spaced instant heart rate readings (in beats per minute) obtained at intervals of 0.5 seconds. Fig. 1 presents the time series plot for T1, showing the heart rate measurements over the studied interval. T3 and T4 were obtained in the same way with 950 and 851 measurements each. Z-score analysis was utilized to identify potential outliers in the dataset. After a thorough investigation, it was determined that these outliers were not due to input errors; therefore, they were retained for further analysis. Min-max normalization was applied to all models to ensure that all values fall within the range of 0 to 1 and to standardize the input data scale across each model. This step is crucial when working with deep learning models because it facilitates faster convergence during training and enhances overall model performance. Normalizing the data helps mitigate the effects of varying feature scales, leading to improved numerical stability and more reliable results.
@@ II-B Framework
For each of the four-time series datasets (T1, T2, T3, T4), individual models were constructed using Seasonal-ARIMA, Prophet, LSTM, TSMixerx, TimesNet, TCN, PatchTST, and iTransformer.
@@ II-C Time Series Models Training
Prophet: We utilized the Facebook Prophet model as another benchmark. This model was employed with parameters such as changepoint_prior_scale, seasonality_prior_scale, and n_changepoints tuned through cross-validation. We also combined with adjustments aimed to optimize forecasting accuracy by controlling the model’s sensitivity to trend changes and seasonal fluctuations. Fig. 2 compares observed data points with the Prophet forecast, highlighting trend adjustments detection.
@@
Long Short-Term Memory (LSTM) Network [9, 10]: The LSTM model included two LSTM layers followed by a dense layer. Heart rate data was normalized using MinMaxScaler. We use three layers of LSTM with hidden dimension of 32. The model was trained using the MSE loss function to enhance performance.
Temporal Convolutional Network (TCN) has been used in time series prediction with huge success. We trained the model using dilated temporal convolutions to capture long-range dependencies in the heart rate data. Five layers of dilated convolutions with residual connections were employed to enhance training efficiency. The heart rate data was normalized before training. We set filter size to be three and dilation factor to be two. The MSE loss function used to guide the training process, ensuring effective learning of temporal patterns.